A pound of tin is worth $1 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $3.30 per pound. How much is a pound of tin worth?

Answer by mananth(16946)

A pound of tin is worth $1 more than a pound of copper. Two pounds of tin are mixed with 8 pounds of copper to make bronze that sells for $3.30 per pound. How much is a pound of tin worth?
let 1 pound of copper cost $ x
Tin will be x+1
..
2(x+1)+8x = 10*3.30
2x+2+8x =33
10x+2=33
-2
10x= 33-2
10x=31
/10
x= 3.1